OpenStudy (anonymous):

Substitute the data provided for speed (v) and starting heights (s) into their proper positions in the formula y = –16x2 + vx + s and write the equation corresponding with each motor in the space provided below. Then identify the a, b, and c values for each where indicated. (6 points total, 2 points each) The equation for motor A is: a = b = c = The equation for motor B is: a = b = c = The equation for motor C is: a = b = c = Motor A speed: 190 Motor B speed: 210 Motor C speed: 230

OpenStudy (anonymous):

I really need help with this and soon so please!

OpenStudy (anonymous):

where are the height values to be substituted ?

OpenStudy (anonymous):

Imagine you are a fountain designer and you need to know how different speed motors will affect the water’s projected height and time in the air. You have various motor speed options but need to choose the one that most closely meets the client’s height request of 690 feet. Your job is to determine which motor speed is the best choice. In all cases, you will use the projectile formula y = –16x2 + vx + s, with starting height of 0 (s = 0).

OpenStudy (anonymous):

Does that help?

OpenStudy (dan815):

just replace the constant variables there with the data

OpenStudy (anonymous):

Can you give me an example, like just do the first one?

OpenStudy (dan815):

Motor A speed: 190

OpenStudy (dan815):

y = –16x2 + 190*x + s

OpenStudy (dan815):

since i dont see any data for the initial displacement u can just say s = 0

OpenStudy (anonymous):

The equation for motor A is: a =-16 b =190 c =0 Does that look good?

OpenStudy (dan815):

ya

OpenStudy (dan815):

y=ax^2+bx+c <---- standard form y = –16x2 + 190x + 0 ^---a ^---b ^----c

OpenStudy (anonymous):

So The equation for motor A is: a =-16 b =190 c =0 The equation for motor B is: a =-16 b =210 c =0 The equation for motor C is: a =-16 b =230 c =0 Like that?

OpenStudy (anonymous):

According to the extra information you have given, 'Y' is supposed to be the final height i.e. 690 feet, and 's' is always zero. Therefore the equation for motor A is: 690=-16x^2+190x+0 => 16x^2 -190x +690 =0 a =16 b =-190 c =690 similarly the others

OpenStudy (dan815):

yep